# Communication and Energy-Aware Multi-UAV Coverage Path Planning for Networked Operations

### v2

@@ Abstract
This paper presents a communication and energy-aware Multi-UAV Coverage Path Planning (mCPP) method for scenarios requiring continuous inter-UAV communication, such as cooperative search and rescue and surveillance missions. Unlike existing mCPP solutions that focus on energy, time, or coverage efficiency, the proposed method generates coverage paths that require minimal communication range to maintain inter-UAV connectivity while also optimizing energy consumption. The mCPP problem is formulated as a multi-objective optimization task, aiming to minimize both the communication range requirement and energy consumption. Our approach significantly reduces the communication range needed for maintaining connectivity while ensuring energy efficiency, outperforming state-of-the-art methods. Its effectiveness is validated through simulations on complex and arbitrary shaped regions of interests, including scenarios with no-fly zones. Additionally, real-world experiment demonstrate its high accuracy, achieving 99% consistency between the estimated and actual communication range required during a multi-UAV coverage mission involving three UAVs.
@@ I INTRODUCTION
In the past few years, the deployment of multiple Unmanned Aerial Vehicles (UAVs) has attracted the interest of many enterprise fields, becoming a powerful tool for professionals to acquire data in a fast and efficient way. Multi-UAV systems have revolutionized various industries by enabling complex missions to be carried out quickly and with precisely, making them invaluable in fields such as surveillance [1], photogrammetry [2], agricultural monitoring, precision farming [3], and Search And Rescue (SAR) [4, 5]. By utilizing multiple UAVs simultaneously, entire Regions Of Interest (ROI) can be surveyed quickly and thoroughly, addressing challenges that were previously difficult to overcome.
The advantages of multi-UAV systems are evident in their ability to rapidly cover large areas and coordinate tasks seamlessly. They efficiently scan vast regions for SAR operations, monitor crop health and manage resources in agriculture [6, 7], and assess disaster-stricken areas by generating detailed maps and 3D models for informed decision-making [8]. These capabilities showcase the transformative impact of multi-UAV systems across diverse sectors, driving innovation and enhancing operational efficiency.
Given the diverse applications of multi-UAV systems, practitioners in the field face significant challenges in maintaining reliable communication networks while executing missions efficiently. In real-world scenarios like coordinated SAR operations, UAVs must continuously share critical information such as victim locations and environmental hazards. These operations are complicated by the limited communication range of onboard modules, which can lead to network fragmentation when UAVs disperse. Traditional star topologies compound this issue by introducing latency and creating a single point of failure, while mesh architectures provide greater reliability through direct peer-to-peer communications and adaptation to dynamic topology changes. These communication requirements directly impact path planning strategies. Consequently, a major challenge in Multi-UAV Coverage Path Planning (mCPP) is generating paths that not only maximize coverage within the ROI and minimize energy consumption, but also maintain communication links throughout the mission. Optimizing all of these factors is an NP-hard problem. Although various solutions have been proposed, they include algorithms that handle No-Fly Zones (NFZs) [9, 10, 11], enhance energy efficiency [10, 12], optimize flight speed [10], and support uneven workload distribution [9, 12].
While research on mCPP has predominantly focused on optimizing energy, coverage, and time, the consideration of communication requirements in coverage path planning remains largely unexplored. In multi-UAV cooperative missions, reliable communication between vehicles is crucial. Designing coverage paths without accounting for communication needs can lead to frequent network disconnections during the mission. This can result in coordination challenges and potential data loss. Maintaining reliable communication is critical in operations such as security surveillance, inspection, and SAR, where it is essential for mission success.
In this paper, we propose a communication and energy aware mCPP algorithm that generates coverage paths with minimal inter-UAV communication range requirements. This ensures that all UAVs remain connected throughout the mission in a mesh network architecture. Simultaneously, the algorithm minimizes energy consumption, maximizes area coverage, and completes the mission efficiently. The algorithm utilizes the groundwork laid by Divide Areas based on Robots initial Positions (DARP) algorithm [9], solving the mCPP problem through area discretization with optimal allocation and constructing a Spanning Tree Coverage(STC) paths on discretized regions. Energy requirement for the coverage paths is estimated using the direct power consumption model from the UAVs.
The primary contribution of this paper is an optimization algorithm that generates coverage paths with minimal inter-UAV communication range requirements. We achieve this by treating the range requirement and energy consumption as hyperparameters in the optimization process, solved using Bayesian optimization techniques across multiple DARP initialization trials. Additionally, we optimize the launch points on generated paths to further minimize communication range requirements. A time-sampling methodology is implemented on these paths, followed by continuous connectivity evaluation to analyze range needs. An optimal shift and rotation technique is applied before discretizing the polygon into cells, enhancing coverage efficiency. This multi-layered optimization framework improves inter-UAV connectivity and energy efficiency, while also supporting complex polygons with NFZs.
@@ II RELATED WORK
Figure 1: Optimization process breakdown: ROI Discretization through optimal shifts and rotation, DARP-MST path generation, time-sampling & energy calculation, launch point optimization by connectivity analysis, Bayesian optimization with range and energy for next DARP initialization
Coverage Path Planning (CPP) is a well-addressed problem, with several solutions already implemented in real-world path planning applications such as floor-cleaning robots, painting systems, UAV survey path planners, and agricultural spraying robots[13, 14, 15]. This section discusses some of the works that have provided key motivation for this paper through their innovative approaches to the problem or optimized results for different flight parameters.
Among the key studies, Comprehensive analyses of CPP techniques are provided in [16, 17], where the authors explore a range of bio-inspired, computational, and analytical CPP methodologies. These studies offer valuable insights into area division and path-planning algorithms for both simple and complex ROIs. They address CPP challenges in both 2D and 3D environments, with and without NFZs, primarily focusing on single-robot applications while also considering multi-robot path planning. The use of trapezoidal cell decomposition [17], STC [9], and energy-aware CPP strategies [10] are highlighted, with emphasis on optimizing flight parameters such as energy consumption, path length, and coverage efficiency. These analyses underscore the complexities of CPP, particularly in cooperative multi-robot scenarios.
Another notable contribution is the introduction of boustrophedon cell decomposition as an alternative to classical trapezoidal decomposition [11, 18]. This technique enables CPP in ROIs with NFZs by partitioning them into cells and using systematic back-and-forth sweeping paths within each cell. A graph traversal method is then used to connect these individual paths into a complete coverage path. Building on this, [19] revisits boustrophedon CPP by replacing the graph traversal method with the Generalized Traveling Salesman Problem (GTSP) [20] to determine the optimal path configuration, minimizing time and directional changes and improving energy efficiency.
@@
All of the discussed CPP solutions utilize area partitioning and optimal allocation. In [25], a formation flying technique was proposed for multi-UAV coverage missions in ROIs without any NFZs. This approach ensures that the inter-UAV distance remains within the range due to the rigid formation shape, thereby maintaining the UAVs within communication range of each other. Another distinct technique involves discretizing the ROI into grid cells and performing STC. In [9], a fast and effective algorithm is proposed by discretizing the ROI into square grid cells and then using a modified Voronoi algorithm to allocate cells to each robot. The STC algorithm [29] is then implemented for each region to generate independent coverage paths. A similar approach is used in [30], where adaptive grid size discretization and the optimal allocation Minimum Spanning Tree (MST) are employed to generate smooth and minimal-cost paths.
To optimize energy in DARP-generated paths, (O-DARP) [12] proposes the use of machine learning techniques for generating energy-efficient coverage paths. In this approach, the energy consumption of paths is mapped to the number of turns, and optimization techniques are applied to generate paths with fewer turns across different DARP initialization.
In contrast to the works discussed above, while all of the proposed approaches present interesting ideas to reduce flight time and energy consumption, they also exhibit several limitations. For instance, many approaches fail to handle arbitrary or complex ROIs, particularly those with NFZs, and do not account for communication constraints that are critical in cooperative missions. Moreover, some methods, like the formation flying technique proposed in [25], focus solely on maintaining inter-UAV distance under range but cannot work for ROIs with NFZs. A comparison of the discussed approaches is summarized in Table I.
@@ V-B Validation in Simulation
The visual representation of five of the generated paths is shown in Fig. 3. The mission altitude used for the Cape shape ROI is 70m above ground lavel and 45 meters for the rest of the ROIs. Additionally, the connectivity estimation for complex and island ROIs is explained in the supplementary video.
@@
Fig. 2 presents the simulation results of the proposed solution across cape, island with NFZs, and rectangular-shaped ROIs, with varying energy-communication trade-off coefficients ($\lambda$). Analyzing the impact of $\lambda$ across different ROIs shows that while the energy-efficient coverage paths have a higher range requirement, the minimum range required paths are not always energy-efficient. In Fig. 2, the Pareto front is marked in orange, indicating the optimal values tuned for different levels of energy and communication awareness. The energy and communication requirements of the EA-mCPP method are shown with a green rectangle mark, while the requirements of the O-DARP algorithm are indicated with a purple triangle mark. The performance of the formation flying strategy is compared in the rectangular ROI in Fig. 2(c).
@@ V-C Real World Experiments
To validate the feasibility of the proposed method in real-world scenarios and assess the accuracy of the simulation, we conducted a multi-UAV cooperative coverage mission with three UAVs. The mission was conducted over the simple polygon ROI mentioned in Table.
@@ VI CONCLUSION
We introduce a novel communication and energy-aware multi-UAV coverage path planning method that combines DARP+STC path generation with iterative connectivity assessment optimization. To the best of our knowledge, our approach is the first to generate communication-aware coverage paths for multi-UAV cooperative missions requiring continuous inter-UAV communication. The method incorporates a communication-energy trade-off parameter, allowing users to balance energy and communication awareness during path generation. Simulations demonstrate that our approach outperforms state-of-the-art methods, reducing communication range requirements by 20-60% and exhibiting better energy awareness compared to existing communication-aware method. The method’s feasibility is validated for arbitrary polygonal shapes and polygons with no-fly zones. Real-world experiments with three UAVs confirm the effectiveness of the proposed solution.

### v3

@@ Abstract
This paper presents a communication and energy-aware multi-UAV Coverage Path Planning (mCPP) method for scenarios requiring continuous inter-UAV communication, such as cooperative search and rescue and surveillance missions. Unlike existing mCPP solutions that focus on energy, time, or coverage efficiency, the proposed method generates coverage paths that minimize a specified combination of energy and inter-UAV connectivity radius. Key features of the proposed algorithm include a simplified and validated energy consumption model, an efficient connectivity radius estimator, and an optimization framework that enables us to search for the optimal paths over irregular and obstacle-rich regions. The effectiveness and utility of the proposed algorithm is validated through simulations on various test regions with and without no-fly-zones. Real-world experiments on a three-UAV system demonstrate the remarkably high 99% match between the estimated and actual communication range requirement.
@@ I Introduction
Coordinated missions performed by teams of Unmanned Aerial Vehicles (UAVs) are becoming increasingly critical for addressing the complex operational challenges encountered in a wide range of scenarios. Multi-UAV systems enable quick completion of complicated missions, such as in surveillance [1], photogrammetry [2], agricultural monitoring, precision farming [3, 4, 5], and Search And Rescue (SAR) [6, 7]. Compared to a single-UAV operation, a multi-UAV system can enhance the mission efficiency and reliability by providing increased coverage area, reduced operational time, improved redundancy, and enhanced collaboration capabilities.
Inter-UAV communication is the cornerstone of many such coordinated missions, e.g. in SAR operations where the UAVs may need to share the locations of victims or environmental hazards, or in surveillance operations requiring cooperative target tracking. For multi-UAV systems that are connected over a mesh network, maintaining inter-UAV communication requires ensuring that the UAVs do not stray too far from each other. In practice, UAVs are generally equipped with small antennas, and therefore the inter-UAV communication range is typically shorter than the UAV-ground control station range. Hence, planning multi-UAV paths without considering inter-UAV communication requirements may result in a fragmented UAV network that is unable to complete the mission.
We consider the Multi-UAV Coverage Path Planning (mCPP) problem, where the goal is to generate multi-UAV paths covering a specified area. Such coverage requirement arises in a number of UAV applications, e.g., photogrammetry, surveillance, environmental monitoring, sensing, and precision farming. The mCPP problem has been widely studied and enhanced algorithms have been proposed to handle No-Fly Zones (NFZs) [8, 9, 10], enhance energy efficiency [9, 11], optimize flight speed [9], and support uneven workload distribution [8, 11]. However, existing works in the literature do not consider the communication requirements.
In this paper, we consider the communication and energy-aware mCPP problem which seeks to find the coverage path that minimizes a combination of energy consumption and connectivity radius. The framework relies on two key building blocks: a light weight energy consumption model and an connectivity radius estimator. The estimates allow us to evaluate candidate paths and search for the optimal path.
To construct candidate paths that support irregular and obstacle-rich regions, we utilize the Divide Areas based on Robots initial Positions (DARP) algorithm [8] for dividing the coverage workload among the given UAVs and the Spanning Tree Coverage (STC) algorithm [12] for generating the paths for each UAV. Subsequently, we utilize Bayesian optimization to find the optimal path that minimizes the specified combination of energy and connectivity radius. In summary, the key contributions of this work are as follows:
  • We propose a parameterizable constant-regime energy model that can be analytically evaluated for any path, achieving less than 5% error in practice.
  • We prove that the mesh-network connectivity radius is Lipschitz continuous, enabling use of the Shubert–Piyavskii global optimizer and yielding significant speedups, especially for larger UAV fleets.
  • We develop a nested Bayesian optimization strategy that efficiently finds near-global optima by pruning candidates using a running median heuristic.
  • The resulting algorithm is a complete, end-to-end planner that supports irregular and obstacle-rich regions of interest (ROIs) with guaranteed coverage.
@@ II RELATED WORK
Figure 1: Optimization process breakdown: ROI Discretization through optimal shifts and rotation, DARP-MST path generation, time-sampling & energy estimation, launch point optimization by connectivity analysis, Bayesian optimization with range and energy for next DARP initialization
Coverage Path Planning (CPP) is a well-studied problem, with several solutions already implemented in real-world path planning applications such as floor-cleaning robots, painting systems, UAV survey path planners, and agricultural spraying robots[13, 14, 15]. This section discusses some of the works that have provided key motivation for this paper through their innovative approaches to the problem or optimized results for different flight parameters.
A number of CPP techniques are reviewed in [16, 17], including bio-inspired, computational, and analytical approaches for both simple and complex ROIs. In this context, it is observed that a number of techniques utilize variants of trapezoidal cell decomposition [17], STC [8], and energy-aware CPP strategies [9], with varying emphasis on optimizing flight parameters such as energy consumption, path length, and coverage efficiency.
Boustrophedon cell decomposition is an alternative to classical trapezoidal decomposition [10, 18] that enables CPP in ROIs with no-fly zones (NFZs) by partitioning them into cells and using systematic back-and-forth sweeping paths within each cell. A graph traversal method is then used to connect these individual paths into a complete coverage path. Building on this, [19] revisits boustrophedon CPP by replacing the graph traversal method with the Generalized Traveling Salesman Problem (GTSP) [20] to determine the optimal path configuration, minimizing time and directional changes and improving energy efficiency.
@@
All of the discussed CPP solutions utilize area partitioning and optimal allocation. In [25], a formation flying technique was proposed for multi-UAV coverage missions in ROIs without any NFZs. This approach ensures that the inter-UAV distance remains within the range due to the rigid formation shape, thereby maintaining the UAVs within communication range of each other. Another distinct technique involves discretizing the ROI into grid cells and performing STC. In [8], a fast and effective algorithm is proposed by discretizing the ROI into square grid cells and then using a modified Voronoi algorithm to allocate cells to each robot. The STC algorithm [12] is then implemented for each region to generate independent coverage paths. A similar approach is used in [29], where adaptive grid size discretization and the optimal allocation Minimum Spanning Tree (MST) are employed to generate smooth and minimal-cost paths. More recently, machine learning techniques have also been used to generate energy-efficient coverage paths [11]. In this approach, the energy consumption of paths is mapped to the number of turns, and optimization techniques are applied to generate paths with fewer turns across different DARP initializations. We remark that many of these techniques cannot handle complex ROIs and none of them consider communication or range requirements. Finally, methods that involve formation flying [25], focus solely on maintaining the inter-UAV distance and cannot work for ROIs with NFZs. A comparison of the discussed approaches is summarized in Table I.
@@ V-B Validation in Simulation
The visual representation of five of the generated paths is shown in Fig. 3. The mission altitude used for the Cape shape ROI is 70m above ground lavel and 45 meters for the rest of the ROIs.
@@
Fig. 2 presents the simulation results of the proposed solution across cape, island with NFZs, and rectangular-shaped ROIs, with varying energy-communication trade-off coefficients ($\lambda$). Analyzing the impact of $\lambda$ across different ROIs shows that while the energy-efficient coverage paths have a higher range requirement, the minimum range required paths are not always energy-efficient. In Fig. 2, the Pareto front is marked in orange, indicating the optimal values tuned for different levels of energy and communication awareness. The energy and communication requirements of the EA-mCPP method are shown with a green rectangle mark, while the requirements of the O-DARP algorithm are indicated with a purple triangle mark. The performance of the formation flying strategy is compared in the rectangular ROI in Fig. 2(c). Additionally, the connectivity estimation for complex and island ROIs is explained in the supplementary video.
@@ V-C Real World Experiments
To validate the feasibility of the proposed method in real-world scenarios and assess the accuracy of the simulation, we conducted a multi-UAV cooperative coverage mission with three UAVs. The mission was conducted over the simple polygon ROI mentioned in Table III.
@@ VI CONCLUSION
We introduce a novel communication and energy-aware multi-UAV coverage path planning method that combines DARP+STC path generation with iterative connectivity assessment optimization. Our approach is the first to generate communication-aware coverage paths for multi-UAV cooperative missions requiring continuous inter-UAV communication. The method incorporates a communication-energy trade-off parameter, allowing users to balance energy and communication awareness during path generation. Simulations demonstrate that our approach outperforms state-of-the-art methods, reducing communication range requirements by 20-60% and exhibiting better energy awareness compared to existing communication-aware method. The method’s feasibility is validated for arbitrary polygonal shapes and polygons with no-fly zones. Real-world experiments with three UAVs confirm the effectiveness of the proposed solution.